# Dynamic Integration of Task-Specific Adapters for Class Incremental Learning

### v1

@@ Introduction
$o$ $=\mathcal{A}(\mathbf{p})=\mathbf{W}^{\top}\mathbf{p}=\sum(\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{\top})^{\top}\mathbf{p}$
$=\sum\mathbf{v}_{i}(\sigma_{i}\mathbf{u}_{i}^{\top}\mathbf{p})=\sum s^{{}^{\prime}}_{i}\mathbf{v}_{i}=\mathbf{V}^{\top}g(\mathbf{p})$ (8)
$g(\mathbf{p})$ $={\text{diag}(\sigma)}^{\top}U^{\top}\mathbf{p}={[s^{{}^{\prime}}_{i}]_{i=1}^{r}}\in\mathbb{R}^{r}.$ (9)
From the above derivation, we can observe that each task adapter allows the model to learn task-specific “keys” and “values” independent of the input features. The output $\mathbf{o}$ for each input token $\mathbf{p}$ is obtained by weighting the task parameter space basis vectors $\mathbf{v}_{i}$ through a linear function $g\left(\cdot;U,\text{diag}(\sigma)\right)$ (a non-linear function when ReLU exists). TSAI ensures that the output $\mathbf{o}$ remains within the task subspace, regardless of the input. We can leverage this property by designing appropriate loss functions to maintain feature consistency under the NECIL setting.
Patch-Level Model Alignment
Due to the limitations of NECIL methods that do not store old class exemplars, mainstream methods do not perform model alignment or only align the classifier.
Based on TSAI’s knowledge retention and reconstruction ability, as well as the rich information conveyed by patch tokens, we propose a patch-level distillation loss to maintain feature consistency and a patch-level old class feature reconstruction method to align the classifier.
@@
As shown in Fig.2 (b), during the training of task $t$, we obtain the output features $\mathbf{X}^{n}$ and $\mathbf{X}^{o}$ for the input image $I$ from both the current model and the model trained on the previous task $t-1$.
$\mathbf{X}^{o}={\left[{\mathbf{p}_{j}^{o}}^{\top}\right]}_{j=0}^{L},\mathbf{X}^{n}={\left[{\mathbf{p}_{j}^{n}}^{\top}\right]}_{j=0}^{L}\in\mathbb{R}^{(L+1)\times d},$ (10)
@@
$\alpha_{\cos}=\frac{\mathbf{p}_{0}^{n}\cdot\mathbf{p}_{j}^{n}}{\|\mathbf{p}_{0}^{n}\|_{2}\|\mathbf{p}_{j}^{n}\|_{2}},\alpha_{\angle}=\frac{\pi-\arccos(-\alpha_{\cos})}{\pi},$ (11)
here, $\alpha_{\cos}$ and $\alpha_{\angle}$ represent the cosine similarity and angular similarity, respectively. We also conduct ablation experiments on different similarity metrics in Table 5.
Secondly, we map patch tokens onto a hypersphere using $L_{2}$ normalization to ensure numerical stability. Then, we measure the feature drifts by calculating the distances between the corresponding tokens on the hypersphere, as follows:
$\mathcal{D}(\mathbf{p}_{j}^{n},\mathbf{p}_{j}^{o})=\|\bar{\mathbf{p}_{j}^{n}}-\bar{\mathbf{p}_{j}^{o}}\|_{2}.$ (12)
We allow greater flexibility for patch tokens that contribute significantly to new tasks. For patches that contribute less to new tasks, we align their tokens with the output of the old model to maintain feature consistency with previous tasks. Thus, the PDL loss function is defined as:
$\mathcal{L}_{pdl}=\frac{1}{L}\sum_{j}\text{SG}\left[\alpha_{\angle}(\mathbf{p}_{0}^{n},\mathbf{p}_{j}^{n})\right]\mathcal{D}(\mathbf{p}_{j}^{n},\mathbf{p}_{j}^{o}),$ (13)
where SG represents the stop gradient operation.
@@
Adam-Prompt-deep 0.28M 36.28B 72.3 78.75 53.46 64.75 86.6 91.42 83.43 89.47
@@
Table 1: Experimental results on four CIL benchmarks with 10 incremental tasks. The best results are marked in bold, and the second-best are underlined. We report the accuracy of compared methods with their source code. Params indicates the trainable parameters in each incremental task. Flops represents the number of floating-point operations required to perform inference on a single image.
Patch-Level Feature Reconstruction: As new incremental tasks are learned, the decision boundaries established from old tasks could be dramatically changed (Zhu et al. 2021b; Wang et al. 2023b). Towards this end, we propose a patch-level feature reconstruction method to generate old class features that adapt to the evolving model, providing better classifier alignment.
Specifically, at each incremental task $t$, we compute and store a class prototype $\bm{\mu}_{k}^{t}$ for each class $k$. When learning new incremental tasks, we apply a normalized distance difference $\delta_{(k,i,j)}$ to measure the relevance between patch token $p_{(i,j)}$ and prototype $\bm{\mu}_{k}^{t}$, where $p_{(i,j)}$ indicates the j-th token of $X_{i}$ within the training batch.
$\delta_{(k,i,j)}=\frac{\alpha_{\text{cos}}(\bm{\mu}_{k}^{t},\mathbf{p}_{(i,j)})-\alpha_{\text{cos}}(\mathbf{p}_{(i,0)},\mathbf{p}_{(i,j)})}{\alpha_{\text{cos}}(\bm{\mu}_{k}^{t},\mathbf{p}_{(i,j)})+\alpha_{\text{cos}}(\mathbf{p}_{(i,0)},\mathbf{p}_{(i,j)})},$ (14)
We retrieve the patch tokens $\mathbf{p}_{(i,j)}$ whose $\hat{\delta}_{(k,i,j)}>0$ and integrate them with prototype $\bm{\mu}_{k}^{t}$ using the exponential moving average (EMA) (Morales-Brotons, Vogels, and Hendrikx 2024) to generate old task feature $\hat{\bm{\mu}}_{k}^{t}$ as follows:
$\hat{\bm{\mu}}_{k}^{t}$ $=\beta\cdot\bm{\mu}_{k}^{t}+(1-\beta)\sum_{i,j}\omega_{(i,j)}\cdot\mathbf{p}_{(i,j)},$ (16)
$\left[\omega_{(i,j)}\right]_{i,j}$ $=\text{softmax}(\left[\hat{\delta}_{(k,i,j)}\right]_{i,j}),\quad\sum_{i,j}\omega_{(i,j)}=1,$ (17)
where $\beta$ is a hyperparameter.
In each training batch, we randomly generate old task features through our PFR and calibrate the classifier using the CrossEntropy loss function as previous methods (Zhang et al. 2023; Zhu et al. 2021b). We provide a more detailed training pipeline in the supplementary material.
Experiments
@@
In this section, we first provide implementation details, then present the experimental results with analyses, and finally show ablation studies and visualization results. More detailed experimental results can be found in the supplementary material.
Implementation Details
@@
Dataset: We conduct experiments on Cifar-100 (Krizhevsky and Hinton 2009), CUB-200 (Wah et al. 2011), ImageNet-R (Hendrycks et al. 2021a), and ImageNet-A (Hendrycks et al. 2021b). These datasets contain typical CIL benchmarks and more challenging datasets with a significant domain gap with ImageNet (i.e., the pre-trained dataset). There are 100 classes in Cifar-100 and 200 classes in CUB, ImageNet-R, and ImageNet-A. For all datasets, we follow the class orders in (Zhou et al. 2024b).
@@
Comparison methods: We compare our method with benchmark PTM-based CIL methods in Table 1 and Table 3. We divide them into three groups: (1) Prompt-based methods: L2P (Wang et al. 2022c), DualPrompt (Wang et al. 2022b), CODA-Prompt (Smith et al. 2023), CPrompt (Gao, Cen, and Chang 2024), ConvPrompt (Roy et al. 2024), Adam-Prompt (Zhou et al. 2024a). (2) Adapter-based methods: Adam-Adapter (Zhou et al. 2024a), EASE (Zhou et al. 2024b), LAE (Gao et al. 2023), (3) Finetuning-based methods: SLCA (Zhang et al. 2023), Adam-Ft (Zhou et al. 2024a).
@@
Training details: We adopt ViT-B/16 (Dosovitskiy et al. 2020) as the pre-trained model, which is pre-trained on ImageNet-21K (Russakovsky et al. 2015). The initial learning rate is set to 0.025 and decays with cosine annealing. We train each task for 20 epochs with a batch size of 32 using Nvidia 3090 GPUs with 24GB of RAM. The down projection of adapters is set to 8 in our main experiments, indicated by DIA-r8 and the hyperparameter $\beta$ is set to 0.7.
@@
Evaluation metric: Following previous papers (Gao et al. 2023; Zhou et al. 2024b), we denote the model’s accuracy after the $t$-th incremental task as $A^{t}$ and use $\bar{A^{t}}=\frac{1}{t}\sum_{i=1}^{t}A^{i}$ to represent the average accuracy over t incremental tasks. For our evaluation, we specifically focus on two key measurements: $A^{T}$ (the accuracy after the final($T$-th) incremental task) and $\bar{A^{T}}$ (the average accuracy across all $T$ incremental tasks).
Comparison with SOTA Methods
@@
In this section, we evaluate our proposed DIA method on the ImageNet-R, ImageNet-A, CUB-200, and Cifar-100 datasets, equally dividing the classes into 10 incremental tasks. As shown in Table 1, our method achieves SOTA average accuracies of 85.61%, 70.43%, 92.13%, and 94.29% on four benchmark datasets, respectively.
@@
Prompt-based Methods: In comparison to prompt-based approaches, the proposed DIA demonstrates comprehensive improvements. When compared with the current SOTA method CPrompt (Gao, Cen, and Chang 2024), our DIA exhibits accuracy enhancements ranging from 1.76% to 4.99% across all datasets. Moreover, the DIA-r08 requires fewer training parameters (merely 0.17M) and reduces the floating-point operations (FLOPS) per image by 24.17%. ConvPrompt (Roy et al. 2024) improves inference efficiency by introducing large language models. Our approach, however, offers advantages in both computational efficiency and accuracy across all datasets, notably achieving 4.06% higher average accuracy on ImageNet-R.
@@
Adapter-based Methods: In comparison to adapter-based methods, our DIA still performs exceptionally well. Utilizing only 14.28% of the trainable parameters needed per incremental task, DIA-r08 outperforms EASE (Zhou et al. 2024b) and Adam (Zhou et al. 2024a) across all benchmark datasets. Moreover, DIA-r08 only needs 17.91B Flops to infer an image, reducing inference consumption by 90% compared to EASE (Zhou et al. 2024b) while improving accuracy. Specifically, DIA-r08 surpasses EASE (Zhou et al. 2024b) by 3.88%, 5.09%, 1.62%, and 1.94% on ImageNet-R, ImageNet-A, CUB-200, and Cifar-100, respectively. Our notable improvements on ImageNet-R and ImageNet-A underscore that incorporating compositionality and leveraging knowledge from old tasks can significantly enhance the learning of new classes that have a domain gap with the pre-trained data.
@@
Ablation Study
DIA PDL PFR ImageNet-R Cifar-100
20.93 40.35 22.17 41.83
77.13 83.87 88.37 92.31
78.18 84.55 89.32 93.48
78.22 84.25 89.81 93.78
79.08 85.61 90.80 94.29
Table 2: Ablation Study on Different Components of the DIA Framework
@@
DIA-r08 (Ours) 76.32 83.51 88.74 93.41
Table 3: Experimental results on four CIL benchmarks with 20 incremental tasks. The best results are marked in bold, and the second-best are underlined.
Component Analysis Our proposed DIA method consists of three main components: 1) Task-Specific Adapter Integration (TSAI), 2) Patch-Level Distillation Loss (PDL), and 3) Patch-Level Feature Reconstruction (PFR). To validate the efficacy of each component, we conduct ablation experiments on two benchmark datasets: Cifar-100, which overlaps with the pre-training data, and ImageNet-R, which exhibits a domain gap from the pre-training data.
The first row in Table 2 shows the average and final accuracy of the pre-trained image encoder with a learnable classifier. The incorporation of the TSAI module yields performance comparable to SOTA methods on both Cifar-100 and ImageNet-R datasets. This indicates that the patch-level integration mechanism atop the adapter effectively reduces interference between new and old tasks while enhancing new task learning, leading to substantial performance gains. The third and fourth rows highlight the importance of model alignment in NECIL. Leveraging TSAI’s retention and reconstruction capabilities, our method gains nearly 1% improvement with the introduction of PDL or PFR. Finally, the fifth row shows that combining compositionality with model alignment achieves SOTA performance, confirming the effectiveness of our framework.
@@
DIA-PFR 79.03 85.62 90.80 94.29
Table 4: Performance comparison with different feature reconstruction (FR) methods
Figure 3: Visualization of different feature reconstruction methods. We use dark colored triangle to represent actual features and light colored circle to represent pseudo features generated by FR (feature reconstruction) methods.
Similarity Metric ImageNet-R Cifar-100
@@
Euclidean $\alpha_{\text{eu}}$ 76.38 81.73 87.42 92.91
Cosine $\alpha_{\text{cos}}$ 78.28 84.76 89.67 93.88
Angular $\alpha_{\angle}$ 79.03 85.61 90.80 94.29
@@
Table 5: Ablation Study on Similarity Metrics in Patch-Level Distillation Loss (PDL)
Long Sequence Incremental Analysis: We evaluate the performance of each method under the condition of long sequences. In this setting, datasets are equally divided into 20 tasks with 10 classes/tasks in ImageNet-R and 5 classes/tasks in Cifar-100, and the results are summarized in Table 3. Our method still maintains excellent performance in terms of $A^{20}$ and $\bar{A}^{20}$. DIA achieves SOTA average accuracies of 83.51% on the ImageNet-R dataset, outperforming the second by 2.05%. We also achieve comparable results with SLCA on Cifar-100, using only 0.17M/86M parameters.
Feature Reconstruction Analysis: We investigate the impact of different old class feature reconstruction methods on classifier alignment. As shown in Table 4, using a Gaussian distribution to simulate old class features yields incremental accuracies of only 93.51% and 93.79% on Cifar-100. In contrast, our proposed PFR method, which leverages TSAI’s capability to retain old knowledge, achieves an average accuracy of 94.29%. This indicates that our PFR-based old feature reconstruction method better adapts to the changes in old class features during the incremental process.
@@
We also visualize the feature distributions of old classes generated by Gaussian sampling and compare them to the actual old class features on new tasks, as shown in Fig.3. In the 10th incremental task, we visualize the first 10 classes. Fig.3 (a) shows features generated using a Gaussian distribution, forming two distinct clusters with actual features, highlighting a noticeable disparity. In contrast, Fig.3 (b) displays features generated by PFR, which closely align with actual features, forming a single cohesive cluster. This confirms the effectiveness of our proposed feature generation method and demonstrates the DIA framework’s ability to preserve and reproduce old knowledge.
@@
PDL ablation: We explore different similarity metrics to assess the contributions of patch tokens to new task learning and analyze their impact on PDL. Specifically, we investigate Euclidean distance ($\alpha_{\text{eu}}$)  (Zhai et al. 2024b), cosine similarity ($\alpha_{\text{cos}}$), and angular similarity ($\alpha_{\angle}$) as shown in Table 5. The experimental results indicate that angular similarity yields the best performance, achieving SOTA results across both datasets. In contrast, Euclidean distance performs the worst, with a significant accuracy drop. We attribute this poor performance to the numerical instability of Euclidean distance, which is highly sensitive to variations in numerical range and more prone to outliers. Additionally, cosine similarity results in a slight performance decline, which occurs because it introduces negative loss values, leading to the offsetting of losses between tokens.
Conclusion

### v2

@@ 4.2 Task-Specific Adapters Integration
$\mathbf{o}$ $=\mathcal{A}(\mathbf{p})=\mathbf{W}^{\top}\mathbf{p}=\sum(\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{\top})^{\top}\mathbf{p}$
$=\sum\mathbf{v}_{i}(\sigma_{i}\mathbf{u}_{i}^{\top}\mathbf{p})=\sum\rho_{i}\mathbf{v}_{i}=\mathbf{V}^{\top}g(\mathbf{p}),$ (9)
$g$ $(\mathbf{p})={\text{diag}(\sigma)}^{\top}\mathbf{U}^{\top}\mathbf{p}={[\rho_{i}]_{i=1}^{r}}\in\mathbb{R}^{r}.$ (10)
From the above derivation, it is evident that each task adapter enables the model to learn task-specific “keys” and “values” independently of the input features. The output $\mathbf{o}$ for each input token $\mathbf{p}$ is obtained by weighting the task parameter space basis vectors $\mathbf{v}_{i}$ through a linear function $g\left(\cdot;\mathbf{U},\text{diag}(\sigma)\right)$, which becomes non-linear in the presence of a ReLU activation. TSAI guarantees that the output discriminative information $\mathbf{o}$ remains within the task subspace, irrespective of the input. This property can be effectively utilized by designing appropriate loss functions to maintain feature consistency under the NECIL setting.
4.3 Patch-Level Model Alignment
Due to the limitations of not storing old class exemplars, mainstream NECIL methods either do not perform model alignment or only align the classifier [49, 45, 7]. Based on TSAI’s knowledge retention and reconstruction capabilities, as well as the rich information conveyed by patch tokens, we propose a patch-level distillation loss to maintain feature consistency and a patch-level old class feature reconstruction method to align the classifier.
@@ 4.3 Patch-Level Model Alignment
As shown in Fig.2 (b), during the training of task $t$, we obtain the output features $\mathbf{X}^{n}$ and $\mathbf{X}^{o}$ from both the current model and the model trained on the previous task $t-1$, each consists of $L+1$ image tokens.
$\mathbf{X}^{o}={\left[{\mathbf{p}_{j}^{t-1}}^{\top}\right]}_{j=0}^{L},\mathbf{X}^{n}={\left[{\mathbf{p}_{j}^{t}}^{\top}\right]}_{j=0}^{L}\in\mathbb{R}^{(L+1)\times d},$ (11)
@@
$\alpha_{\cos}=\frac{\mathbf{p}_{0}^{t}\cdot\mathbf{p}_{j}^{t}}{\|\mathbf{p}_{0}^{t}\|_{2}\|\mathbf{p}_{j}^{t}\|_{2}},\alpha_{\angle}=\frac{\pi-\arccos(-\alpha_{\cos})}{\pi},$ (12)
where $\alpha_{\cos}$ and $\alpha_{\angle}$ represent the cosine similarity and angular similarity, respectively. We also conduct ablation experiments on different similarity metrics in Table 5.
Afterward, we map patch tokens onto a hypersphere using $L_{2}$ normalization to ensure numerical stability. We measure the feature drifts $\mathcal{D}$ by calculating the distances between the corresponding tokens on the hypersphere, as follows:
$\mathcal{D}(\mathbf{p}_{j}^{t},\mathbf{p}_{j}^{t-1})=\|\bar{\mathbf{p}}_{j}^{t}-\bar{\mathbf{p}}_{j}^{t-1}\|_{2}.$ (13)
We allow greater flexibility for patch tokens that contribute significantly to new tasks. For patches that contribute less to new tasks, we align their tokens with the output of the old model to maintain feature consistency with previous tasks. Finally, the PDL loss is defined as:
$\mathcal{L}_{pdl}=\frac{1}{L}\sum_{j=1}^{L}\alpha_{\angle}(\mathbf{p}_{0}^{t},\mathbf{p}_{j}^{t})\mathcal{D}(\mathbf{p}_{j}^{t},\mathbf{p}_{j}^{t-1}).$ (14)
Patch-Level Feature Reconstruction: As new incremental tasks are learned, the decision boundaries established from old tasks may undergo significant changes [51, 37]. Towards this end, we propose a patch-level feature reconstruction method to generate old class features that adapt to the evolving model, providing better classifier alignment.
Specifically, at each incremental task $t-1$, we compute and store a class prototype $\bm{\mu}_{k}^{t-1}$ for each class $k$. When learning new incremental tasks, we apply a relative similarity difference $\delta_{(k,i,j)}$ to measure the relevance between patch token $p_{(i,j)}$ and prototype $\bm{\mu}_{k}^{t-1}$, where $p_{(i,j)}$ indicates the $j$-th token of $X_{i}$ within the training batch.
$\delta_{(k,i,j)}=\frac{\alpha_{\text{cos}}(\bm{\mu}_{k}^{t-1},\mathbf{p}_{(i,j)})-\alpha_{\text{cos}}(\mathbf{p}_{(i,0)},\mathbf{p}_{(i,j)})}{\alpha_{\text{cos}}(\bm{\mu}_{k}^{t-1},\mathbf{p}_{(i,j)})+\alpha_{\text{cos}}(\mathbf{p}_{(i,0)},\mathbf{p}_{(i,j)})},$ (15)
where $\alpha_{cos}$ represents the cosine similarity.
Afterward, we retrieve the patch tokens $\mathbf{p}_{(i,j)}$ whose $\hat{\delta}_{(k,i,j)}>0$ and integrate them with prototype $\bm{\mu}_{k}^{t-1}$ using the convex combination to generate old task feature $\hat{\bm{\mu}}_{k}^{t-1}$ as follows:
$\hat{\bm{\mu}}_{k}^{t-1}=\beta\cdot\bm{\mu}_{k}^{t-1}+(1-\beta)\sum_{i,j}\omega_{(i,j)}\cdot\mathbf{p}_{(i,j)},$ (17)
$\left[\omega_{(i,j)}\right]_{i,j}=\text{softmax}(\left[\hat{\delta}_{(k,i,j)}\right]_{i,j}),$ (18)
where $\beta$ is a hyperparameter.
In each training batch, we randomly generate old task features through our PFR and align the classifier through the cross-entropy loss $\mathcal{L}_{CE}$.
4.4 Optimization Objective
We structure the optimization process into two distinct stages: new task learning and classifier alignment. To ensure feature consistency, the objective function for new task learning is defined as follows:
$\mathcal{L}_{obj}=\mathcal{L}_{CE}+\lambda\mathcal{L}_{pdl},$ (19)
where $\lambda$ is a hyperparameter to balance the contributions of the two losses.
To further refine the classifier, classifier alignment is executed after new task learning, following the pipeline detailed in [45, 51]. Specifically, during the classifier alignment phase for task $t$, we randomly select $N$ class prototypes (set to 32 in our implementation), denoted as $\{\mu_{k}\}_{i=1}^{N},k\in C^{1:t}$, within each training batch and generate pseudo-features $\{\hat{\mu}_{k}\}_{i=1}^{N}$ via the PFR method. These pseudo-features, together with the training samples, are then used as inputs to the classifier, which is subsequently fine-tuned with the cross-entropy loss $\mathcal{L}_{CE}$. We provide a more detailed training pipeline in the supplementary material.
5 Experiments
@@
Adam-Prompt-deep 0.28M 36.28B 72.30 78.75 53.46 64.75 86.6 91.42 83.43 89.47
@@
Table 1: Experimental results on four CIL benchmarks with 10 incremental tasks. The best results are marked in bold, and the second-best are underlined. We report the accuracy of compared methods with their source code. Params indicates the average trainable parameters in each incremental task. Flops represents the number of floating-point operations required to perform inference on a single image.
@@
In this section, we first provide implementation details, then present the experimental results with analyses, and finally show ablation studies and visualization results.
5.1 Implementation Details
@@
Dataset: We conduct experiments on Cifar-100 [17], CUB-200 [34], ImageNet-R [9], and ImageNet-A [10]. These datasets contain typical CIL benchmarks and more challenging datasets with a significant domain gap with ImageNet (i.e., the pre-trained dataset). There are 100 classes in Cifar-100 and 200 classes in CUB-200, ImageNet-R, and ImageNet-A. For all datasets, we follow the class orders in [49].
@@
Comparison methods: We compare our method with benchmark PTM-based CIL methods in Table 1 and Table 2. These methods are categorized into three groups: (1) Prompt-based methods: L2P [39], DualPrompt [38], CODA-Prompt [33], CPrompt [7], ConvPrompt [31], and Adam-Prompt [47] (2) Adapter-based methods: Adam-Adapter [47], EASE [49], LAE [5], InfLoRA [25], and C-ADA [6]. (3) Finetuning-based methods: SLCA [45] and Adam-Ft [47].
@@
Training details: We adopt ViT-B/16 [4] as the pre-trained model, which is pre-trained on ImageNet-21K [32]. The initial learning rate is set to 0.015 and decays with cosine annealing. We train each task for 20 epochs with a batch size of 32 using Nvidia 3090 GPUs with 24GB of RAM. The down projection of adapters is set to 8 in our main experiments. The hyperparameter $\beta$ and $\lambda$ are set to 0.7 and 0.1, respectively.
@@
Evaluation metric: Following previous papers [5, 49], we denote the model’s accuracy after the $t$-th incremental task as $A^{t}$ and use $\bar{A}^{t}=\frac{1}{t}\sum_{i=1}^{t}{A^{i}}$ to represent the average accuracy over t incremental tasks. For our evaluation, we specifically focus on two key measurements: $A^{T}$ (the accuracy after the final ($T$-th) incremental task) and $\bar{A}^{T}$ (the average accuracy across all $T$ incremental tasks).
5.2 Comparison with SOTA Methods
@@
In this section, we evaluate our proposed DIA method on the ImageNet-R, ImageNet-A, CUB-200, and Cifar-100 datasets, equally dividing the classes into 10 incremental tasks. As shown in Table 1, our method achieves SOTA average accuracies of 85.61%, 71.58%, 93.21%, and 94.29% on four benchmark datasets, respectively.
@@
Prompt-based Methods: In comparison to prompt-based approaches, the proposed DIA demonstrates comprehensive improvements. When compared with the current SOTA method CPrompt [7], our DIA exhibits accuracy enhancements ranging from 1.76% to 4.99% across all datasets. Moreover, the DIA requires fewer training parameters (merely 0.17M) and reduces the floating-point operations (FLOPS) per image by 24.17%. ConvPrompt [31] improves inference efficiency by introducing large language models. Our approach, however, offers advantages in both computational efficiency and accuracy across all datasets, notably achieving 4.06% higher average accuracy on ImageNet-R.
@@
Adapter-based Methods: In comparison to adapter-based methods, our DIA still performs exceptionally well. Utilizing only 14.28% of the trainable parameters needed per incremental task, DIA outperforms EASE [49] and Adam [47] across all benchmark datasets. Moreover, DIA only needs 17.91B Flops to infer an image, reducing inference consumption by 90% compared to EASE [49] while improving accuracy. Specifically, DIA surpasses EASE [49] by 3.88%, 6.24%, 1.62%, and 1.94% on ImageNet-R, ImageNet-A, CUB-200, and Cifar-100, respectively. Our notable improvements on ImageNet-R and ImageNet-A underscore that incorporating compositionality and leveraging knowledge from old tasks can significantly enhance the learning of new classes that have a domain gap with the pre-trained data.
@@
5.3 Ablation Study
@@
InfLoRA 71.01 77.28 85.80 91.51
DIA (Ours) 76.32 83.51 88.74 93.41
Table 2: Experimental results with 20 incremental tasks. The best results are marked in bold, and the second-best are underlined.
DIA PDL PFR ImageNet-R Cifar-100
20.93 40.35 22.17 41.83
77.13 83.87 88.37 92.31
78.18 84.55 89.32 93.48
78.22 84.25 89.81 93.78
79.03 85.61 90.80 94.29
Table 3: Ablation Study on Different Components of the DIA Framework.
Long Sequence Incremental Analysis: We evaluate the performance of each method under the long sequence setting where datasets are equally divided into 20 tasks with 10 classes/tasks in ImageNet-R and 5 classes/tasks in Cifar-100. As shown in Table 2, our method still maintains excellent performance in terms of $A^{20}$ and $\bar{A}^{20}$. Our DIA achieves SOTA average accuracies of 83.51% on the ImageNet-R dataset, outperforming the second by 2.05%. Our DIA also achieve comparable results with SLCA on Cifar-100, using only 0.17M/86M parameters.
Component Analysis: Our DIA comprises three main components: 1) Task-Specific Adapter Integration (TSAI), 2) Patch-Level Distillation Loss (PDL), and 3) Patch-Level Feature Reconstruction (PFR). To assess the impact of each component, we perform ablation studies on Cifar-100 and ImageNet-R.
The first row in Table 3 reports the average and final accuracy of the pre-trained image encoder with a learnable classifier. Adding the TSAI module achieves performance comparable to SOTA on both datasets, showing that patch-level integration reduces task interference and improves new task learning. The third and fourth rows further demonstrate the importance of patch-level model alignment in NECIL, with nearly 1% gains from PDL or PFR integration. Finally, the fifth row confirms that combining compositionality and model alignment yields SOTA results, validating our framework’s effectiveness.
@@
DIA-PFR 79.03 85.61 90.80 94.29
Table 4: Performance comparison with different feature reconstruction (FR) methods.
Figure 3: Visualization of different feature reconstruction methods. Dark-colored triangles represent actual features, while light-colored circles denote pseudo features generated by feature reconstruction (FR) methods. The decision boundaries formed by these pseudo features are also visualized through the background color. $RE^{Gau}$ and $RE^{PFR}$ indicate regions with inaccurate decision boundaries. The features reconstructed by PFR closely align with the actual features of old classes, resulting in more accurate decision boundaries.
Distillation Loss ImageNet-R Cifar-100
@@
$\mathcal{L}_{pdl}$ w/ $\alpha_{\text{eu}}$ 76.38 81.73 87.42 92.91
$\mathcal{L}_{pdl}$ w/ $\alpha_{\text{cos}}$ 78.28 85.16 89.67 93.88
$\mathcal{L}_{pdl}$ w/ $\alpha_{\angle}$ (ours) 79.03 85.61 90.80 94.29
$\mathcal{L}_{fd}$ 76.67 83.82 88.31 92.97
$\mathcal{L}_{pdl}$ w/ cls 77.01 83.87 88.87 93.02
$\mathcal{L}_{pdl}$ w/o cls (ours) 79.03 85.61 90.80 94.29
@@
Table 5: Ablation Study on Similarity Metrics in $\mathcal{L}_{pdl}$ and other distillation methods.
Figure 4: Ablation experiments on hyperparameters $\beta$ and $\lambda$ conducted on the CIFAR-100 dataset.
Feature Reconstruction Analysis: We investigate the impact of different pseudo-feature reconstruction methods on classifier alignment. As shown in Table 4, constructing old class features using a Gaussian distribution results in accuracies of 83.72% and 93.51% on ImageNet-R and CIFAR-100, respectively. In contrast, our proposed PFR method leverages rich semantics from patch tokens to achieve average accuracies of 85.61% and 94.29%, outperforming other approaches. This demonstrates the effectiveness of PFR in aligning the classifier with old-class features, leading to more accurate decision boundaries.
@@
We also visualize the pseudo-feature distributions generated by Gaussian sampling and compare them to the actual old class features on new tasks, as shown in Fig.3. After learning the last incremental task, we visualize the feature distribution of the first 20 classes. Fig.3 (a) shows pseudo features generated using a Gaussian distribution, forming two distinct clusters with actual features, highlighting a noticeable disparity. In contrast, Fig.3 (b) displays pseudo features generated by PFR, which closely align with actual features, forming a single cohesive cluster. This further results in a more clearly distinguishable decision boundary, as depicted in Fig.3 (b) and (c), thereby confirming the effectiveness of our proposed feature reconstruction method.
@@
PDL Ablation: We explore different similarity metrics to evaluate patch token contributions to new task learning and their impacts on PDL. Specifically, we investigate Euclidean distance ($\alpha_{\text{eu}}$[43], cosine similarity ($\alpha_{\text{cos}}$), and angular similarity ($\alpha_{\angle}$) as shown in Table 5. The experimental results indicate angular similarity outperforms others, achieving SOTA results on both datasets. In contrast, Euclidean distance performs the worst, with a significant accuracy drop. We attribute this poor performance to the numerical instability of the Euclidean distance, which is more sensitive to variations in the numerical range and more prone to outliers.
Additionally, ablation experiments on traditional distillation loss are presented. In Table 5, $\mathcal{L}_{fd}$ [52, 51, 37] refers to applying $L1$ constraints on all image tokens, while $\mathcal{L}_{pdl}$ w/ cls [43] applies the PDL constraint to both the class and patch tokens. The results show that our proposed PDL offers a better trade-off between stability (regularization of feature shift) and plasticity (ability to learn new tasks).
Hyperparameter Ablation We conduct ablation studies for the hyperparameters $\beta$ and $\lambda$ in Fig.4. As depicted in the figure, larger $\lambda$ which results in greater regularization hinders the learning process. Small $\beta$ may lead to inaccurate classifier alignment. Therefore, we suggest the optimal hyperparameter combination of $\beta=0.7$ and $\lambda=0.1$.
6 Conclusion